OpenStudy (anonymous):

"Consider w = z/(z^2+1), where z=x+iy, y \neq 0 and {z^2+1} \neq 0. Given that Im (w) = 0, show that | z | = 1." I would appreciate it if someone pointed me towards the right direction and showed one or two of the first steps, but preferably not the entire answer.

OpenStudy (anonymous):

Written out again because it looks all messed up: \[\frac{ z }{ z^2+1 } , z=x+iy , y \neq 0 , z^2+1 \neq 0 \]

OpenStudy (loser66):

@terenzreignz

terenzreignz (terenzreignz):

I usually decompose the entire thing... how about expressing w as \[\Large \frac{x+\color{blue}iy}{(x+\color{blue}iy)^2 + 1}\] and separate the real and imaginary parts?

OpenStudy (anonymous):

so you mean expanding the brackets?

OpenStudy (anonymous):

\[\frac{ x+iy }{ x^2+2xyi-y^2+1 }\]

terenzreignz (terenzreignz):

Yeah, that's it, and then, well, the real part and the imaginary parts aren't quite separated yet... \[\large \frac{x+\color{blue}iy}{x^2 -y^2 +1 + 2xy\color{blue}i}\] Let's try multiplying the numerator and denominator by the conjugate of the denominator?

terenzreignz (terenzreignz):

Mind you, I don't really know what I'm doing, I'm just hoping that something might come out :3

OpenStudy (anonymous):

hahah alright then :D

OpenStudy (anonymous):

so the conjugate would be \[x^2-y^2+1-2xyi\] ?

terenzreignz (terenzreignz):

That's right...

OpenStudy (anonymous):

So it gives: \[\frac{(x+iy)(x^2-y^2+1-2xyi) }{(x^2-y^2+1+2xyi) (x^2-y^2+1-2xyi) }\] I'm not sure how to proceed as the powers will get pretty big if we multiply it out

terenzreignz (terenzreignz):

Don't mind the denominator, that's sure to be some real number anyway :) Focus on the numerator, multiply the polynomials, and separate the real and the imaginary part...

OpenStudy (anonymous):

so:\[x^3-y^2x+x-x^2yi-y^3i+yi+2xy^2\]

OpenStudy (anonymous):

But to know Im (w) don't we need the denominator also?

terenzreignz (terenzreignz):

Ah, I see now... the magic...

terenzreignz (terenzreignz):

Here, we have this... \[\Large \frac{x^3-y^2x+x+2xy^2-x^2yi-y^3i+yi }{(x^2-y^2+1+2xyi) (x^2-y^2+1-2xyi) }\]

OpenStudy (anonymous):

I think the 2xy^2 should be 2xy^2i

terenzreignz (terenzreignz):

No it shouldn't, because it was the result of you multiplying \(\large y\color{blue}i\) by \(\large -2xy\color{blue}i\)

terenzreignz (terenzreignz):

So, ready to proceed? ^_^

OpenStudy (anonymous):

Aaa I see

terenzreignz (terenzreignz):

Shall we?

OpenStudy (anonymous):

Yep :)

terenzreignz (terenzreignz):

So, the denominator is simply \(z^2 + 1\) multiplied by its conjugate, yes? so, it becomes just \(\large |z^2 +1|^2\) a real number, which isn't zero since \(|z^2+1|\ne0\)

terenzreignz (terenzreignz):

You following me so far?

OpenStudy (anonymous):

yes that seems to be correct

OpenStudy (anonymous):

you make it pretty clear so it's easy to follow :D

terenzreignz (terenzreignz):

So, the entire thingy is just\[\Large \frac{x^3-y^2x+x+2xy^2-x^2yi-y^3i+yi }{|z^2+1|^2 }\] I actually just did that to lessen the scary-factor of the denominator :3

OpenStudy (asnaseer):

there is a much simpler way of solving this

terenzreignz (terenzreignz):

I will defer :) Complex analysis is not my forte... but maybe you guys could just let me finish this? Pretty please? ^_^

OpenStudy (asnaseer):

I am taking account of the fact that:\[\log{w}=0\implies\log{\frac{z}{z^2+1}}=0\implies z=z^2+1\]

OpenStudy (anonymous):

Could you explain that part?

terenzreignz (terenzreignz):

log? How does log factor into all of this? D:

OpenStudy (asnaseer):

if \(\log(x)=0\implies x=e^0=1\)

OpenStudy (asnaseer):

aren't we given \(\log(w)=0\)

terenzreignz (terenzreignz):

No... we're given Im(w) = 0 D:

OpenStudy (anonymous):

I don't think we're given that, although I could be mistaken

OpenStudy (asnaseer):

I thought 'lm' was 'ln' mistyped :)

OpenStudy (anonymous):

:D Would you happen to have other ideas or should terenzreignz continue from where we left off?

OpenStudy (asnaseer):

strangely enough if it was 'ln' then it leads to the same answer! :)

OpenStudy (anonymous):

Yes it does, I just realized that :D Maybe there's something to it....

terenzreignz (terenzreignz):

It's Terence (or TJ) @Teinis >.> Back to reality, I suppose :) We may now neatly separate the real part from the imaginary part, like this... \[\Large \frac{\color{red}{x^3-y^2x+x+2xy^2}+\color{blue}{(-x^2y-y^3+y)}\color{green}i }{|z^2+1|^2 }\]

terenzreignz (terenzreignz):

So, in effect, \[\Large Im(w) = \frac{-x^2y-y^3+y}{|z^2+1|^2}=0\]

terenzreignz (terenzreignz):

Catch me so far?

OpenStudy (anonymous):

Yep sounds good

OpenStudy (asnaseer):

@terenzreignz has it spot on! sorry for my mis understanding of your question. :)

terenzreignz (terenzreignz):

Well, we have established that \(\large |z^2 + 1|^2\) is a nonzero real number, so it may be multiplied to both sides of the equation to give... (and cancel out) \[\Large -x^2y-y^3+y = 0\] aye?

OpenStudy (anonymous):

aaaa Im starting to see where you're going with this

OpenStudy (anonymous):

Yep that seems good

terenzreignz (terenzreignz):

You want to continue it?

OpenStudy (anonymous):

I could give it a try :)

terenzreignz (terenzreignz):

Hint: Factor ;)

OpenStudy (anonymous):

We can factorize it as \[y(-x^2-y^2+1) = 0\]

terenzreignz (terenzreignz):

mhmm <nods>

OpenStudy (anonymous):

and since we know: \[y \neq 0, that means -x^2-y^2+1 = 0\]

terenzreignz (terenzreignz):

yes...

terenzreignz (terenzreignz):

Stuck?

OpenStudy (anonymous):

Actually 'I got stuck again, not quite sure how to continue now...

OpenStudy (anonymous):

NO WAIT!

terenzreignz (terenzreignz):

waiting...

OpenStudy (anonymous):

Omg I just had a light bulb moment

terenzreignz (terenzreignz):

Do elaborate ;)

OpenStudy (anonymous):

\[-x^2-y^2+1 = x^2+y^2-1\]

terenzreignz (terenzreignz):

huh? That makes no sense...

OpenStudy (anonymous):

I meant if you multiply both sides by -1 you can get it into that form

terenzreignz (terenzreignz):

^THAT makes sense

OpenStudy (anonymous):

\[x^2+y^2 = 1\]

terenzreignz (terenzreignz):

<drum roll>

OpenStudy (anonymous):

and since \[\left| z \right| = \sqrt{x^2+y^2}\]

terenzreignz (terenzreignz):

<drum roll gets faster>

OpenStudy (anonymous):

\[\sqrt{x^2+y^2} = 1\]

OpenStudy (anonymous):

so \[\left| z \right| = 1\]

OpenStudy (anonymous):

!!!

terenzreignz (terenzreignz):

<applause>

terenzreignz (terenzreignz):

Heh... not bad for a kid that didn't know what he was doing just about a half-hour ago, huh? ^_^

OpenStudy (anonymous):

I would like to thank my mother and father.....

OpenStudy (anonymous):

and tj ;)

terenzreignz (terenzreignz):

hehe :D

OpenStudy (anonymous):

yeah I guess, now that I think about it it doesn't seem that hard, you just need someone to walk you through some of the process

OpenStudy (anonymous):

Thanks a lot for helping me with that :)

terenzreignz (terenzreignz):

Yes... and I didn't really tell you the answer, now did I? ;)

OpenStudy (anonymous):

hahah nope ;)

terenzreignz (terenzreignz):

You actually did most of the dirty work yourself... I detest algebra XD

OpenStudy (anonymous):

yeah those letters can be quite messy

OpenStudy (anonymous):

Well, I guess it's time to wrap this question up, thanks again!

terenzreignz (terenzreignz):

Signing off now. --------------------------- Terence out